OpenStudy (amtran_bus):

Starting with the graph of y=e^x write the equation of the graph that results from reflecting about the line y = 7.

OpenStudy (amtran_bus):

Starting with the graph of y=e^x write the equation of the graph that results from reflecting about the line y = 7. Is it either y= 14-e^x OR y=e^x-14?

OpenStudy (amtran_bus):

Choices: see attachment

OpenStudy (irishboy123):

i think you are right first time the x axis reflects to y = 14. for very small x your new function will be 14 the y intercept reflects to y = 13. but check this fpr yourself.....

OpenStudy (amtran_bus):

@IrishBoy123 thanks I thought that was right.

undeadknight26 (undeadknight26):

Why so many people on one question? Go help other people. @IrishBoy123 has this.

OpenStudy (amtran_bus):

Will do.

OpenStudy (amtran_bus):

Yes, thanks irishboy